Cheyenne Co. is building a new hockey arena at a cost of $2,460,000. It received a downpayment of $500,000 from local businesses to support the project, and now needs to borrow $1,960,000 to complete the project. It therefore decides to issue $1,960,000 of 10%, 10-year bonds. These bonds were issued on January 1, 2019, and pay interest annually on each January 1. The bonds yield 9%.

(a)

Prepare the journal entry to record the issuance of the bonds on January 1, 2019. (Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to 0 decimal places e.g. 58,971. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date Account Titles and Explanation Debit Credit

January 1, 2019
Answer #1
Proceeds from bond is calculated as present value of coupon amount and present value of par value of bond
Proceeds from bond 1960000*10%*PV of annuity of $1(i=9%,n=10) + 1960000*PV of $1 (i=9%, n=10)
Proceeds from bond (196000*6.41766)+(1960000*0.42241)
Proceeds from bond $2,085,786
Record journal entry for issue of bonds
General Journal Debit Credit
Cash $2,085,786
Bonds payable $1,960,000
Premium on bonds payable (2085786-1960000) $125,786